How many combinations are there for getting exactly 3 tails in 8 fair coin flips? What is the probability that you will land on exactly 3 tails in 8 fair coin flips?
Answer #1

3 tails caa be obtained from 8 fair coin flips by,

8C3 = 56 ways

Therefore there are 56 combinations.

The probability that you will land on exactly 3 tails in 8 fair coin flips is,

56 * (1/2)8 = 56 * (0.5)8 = 0.21875

Probability = 0.21875